OpenStudy (anonymous):

If a,b,c are in A.P then show that :- a^2(b+c) , b^2(c+a) , c^2(a+b) are in A.P

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

The most direct way is to assume a=a b=a+d c=a+2d since a,b,c are in AP. substitute the above into the three expressions, E1,E2,E3 respectively. Calculate and check if E3-E2=E2-E1. If so, the three expressions are in AP. I get E3-E2=E2-E1=2d^3+6ad^2+3a^2d.

OpenStudy (anonymous):

But how to actually prove it

OpenStudy (mathmate):

Without loss of generality, you can also assume a=0, which simplifies things a lot.

OpenStudy (mathmate):

If a^2(b+c) , b^2(c+a) , c^2(a+b) are in A.P, then they should have a common difference. All you need to prove is that if a,b,c are in AP, then c^2(a+b)-b^2(c+a) = b^2(c+a) - a^2(b+c) =D = common difference, hence the three expressions are in AP.

OpenStudy (mathmate):

For details, see initial post above.